• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Folts

    Log entry time 13:21:39 on March 17, 2012

    Entry number 367138


    Call in for a VESDA alarm in Hall A

    Noted that the sensors in the dome and south wall were highest and the the septum had tripped off

    Entered the Hall after a full sweep in controlled access after looking at the radiation map

    Went to the big box power supply to start with and noted that the power was off.

    Called Jack

    Looked at the HKS and noted that it was faulted

    Went to the switch gear to check the CB for Big Box and found it tripped.

    Called MCC to shut down FZ magnets

    Dave Kausch entered the Hall.

    Jack and Tim Minga enter Hall

    Left Hall to go to restricted access.

    Target group locked out target power supply and I went to get the key to the security gates

    While I was gone the others noted soot at the top of septum

    Locked out all power supplies

    Opened gates and found damage to the power connection on the top of the right side of septum.

    See Jacks write up

    Informed Run coordinator and Hall leader of status and recommended that we hold further operations until Monday for Radcon clearance and a thorough investigation.

    Re-locked security gates